### v1

# Hypothesis Generation with Large Language Models

Abstract
Effective generation of novel hypotheses is instrumental to scientific progress. So far, researchers have been the main powerhouse behind hypothesis generation by painstaking data analysis and thinking (also known as the Eureka moment). In this paper, we examine the potential of large language models (LLMs) to generate hypotheses. We focus on hypothesis generation based on data (i.e., labeled examples). To enable LLMs to handle arbitrarily long contexts, we generate initial hypotheses from a small number of examples and then update them iteratively to improve the quality of hypotheses. Inspired by multi-armed bandits, we design a reward function to inform the exploitation-exploration tradeoff in the update process. Our algorithm is able to generate hypotheses that enable much better predictive performance than few-shot prompting in classification tasks, improving accuracy by 31.7% on a synthetic dataset and by 13.9%, 3.3% and, 24.9% on three real-world datasets. We also outperform supervised learning by 12.8% and 11.2% on two challenging real-world datasets. Furthermore, we find that the generated hypotheses not only corroborate human-verified theories but also uncover new insights for the tasks.
1 Introduction
Hypothesis generation drives scientific progress. Mendel’s hypothesis on allele pairs lays the foundation for modern genetics; Einstein’s hypothesis in general theory of relativity led to the prediction and subsequent confirmation of gravitational waves. In the context of language modeling, the hypothesis on scaling law inspires recent progress in large language models (LLMs) (Chowdhery et al., 2022). Despite the importance of hypothesis generation, as Ludwig & Mullainathan (2024) points out, science has been curiously asymmetric. While many scientific publications present extensive formal and empirical evaluation of hypotheses, the generation of hypotheses happens off-stage by researchers. In order to generate novel hypotheses, researchers may read literature, analyze data, pick the brain of each other, and even “hallucinate” (see Kekulé’s discovery of the structure of the benzene molecule (Rothenberg, 1995)). Given the rise of large language models (Brown et al., 2020; Anthropic, 2023; OpenAI, 2023b), we examine their potential of providing much needed assistance in hypothesis generation in this work.
In particular, we focus on hypothesis generation based on data, a common approach in empirical sciences. Our main question is how we can enable LLMs to generate hypotheses of high-quality. While one can easily prompt LLMs to generate hypotheses, LLMs may not be able to effectively leverage the input examples in a single long prompt. Moreover, it is important to have measures of quality in the generation process so that we can filter bad hypotheses and come up with better ones. These two observations motivate us to start with a setup analogous to supervised learning. We can iteratively prompt an LLM to generate hypotheses based on the training examples and use training accuracy as a measure of quality to guide the generation process. Conveniently, we can also evaluate the quality of the final generated hypotheses with their performance on held-out examples, similar to supervised learning.
To generate high-quality hypotheses with LLMs, we propose an algorithm inspired by the upper confidence bound algorithm in multi-armed bandits (Auer, 2002) (HypoGeniC, Hypothesis Generation in Context; see Figure 1). Given initial hypotheses generated from a small number of examples, we need to assess their quality and propose new hypotheses to address their deficiencies. To navigate this exploration-exploitation tradeoff, we introduce a reward function and evaluate the top $k$ hypotheses for each training example. We maintain a wrong example bank to capture the gap in knowledge of the current hypotheses pool, and generate new hypotheses based on the wrong example bank to close the gap.
Figure 1: Illustration of HypoGeniC. During update stage, we evaluate the top $k$ hypotheses on each new training example and update the reward based on the prediction corretness. If the number of hypotheses that got the example wrong exceeds a certain threshold, we add the example to a wrong example bank. The wrong example bank is then used to generate new hypotheses.
The generated hypotheses naturally enable an interpretable hypothesis-based classifier. We propose a suite of inference strategies given a set of hypotheses. We apply our method to one synthetic task where there is a single known valid hypothesis and three real-world tasks (Deceptive reviews, and Tweet popularity). The real-world tasks focus on deception detection and message popularity prediction, which are known to be challenging even for humans (Ott et al., 2011; Salganik et al., 2006). Our algorithm can recover the hypothesis in the synthetic task and also provide useful hypotheses for the real-world tasks. In fact, our generated hypotheses consistently outperform few-shot in-context learning baselines across all four tasks (31.7% in Shoe sales, 13.9% in Deceptive reviews, 3.3% in and 24.9% in Tweet popularity). The predictive performance matches and even outperforms oracle supervised learning with RoBERTa except in Deceptive reviews.
It is important to emphasize that although the utility of hypotheses in assisting downstream classification serves as an indicator for LLMs’ ability to generate hypotheses, our primary interest lies in the quality of the hypotheses. Thus, it is critical for the hypotheses to be interpretable beyond the LLM used to produce the hypotheses. We show that hypotheses generated by one LLM (e.g., GPT-3.5-turbo) can be used to make accurate inference by another LLM (e.g., Mixtral). On an out-of-distribution dataset for Deceptive reviews, we can even outperform the oracle fine-tuned RoBERTa. Such cross generalization provides strong evidence that we are able to generate hypotheses of high quality. Furthermore, through a qualitative analysis, our generated hypotheses not only confirm theories from existing literature but also provide new insights about the task. For instance, one novel hypothesis is that “reviews that mention personal experiences or special occasions, such as birthdays, anniversaries, or weddings, are more likely to be truthful”. We encourage future research on deception detection to explore these novel hypotheses.
Our work is connected to many recent studies on using LLMs to propose “hypotheses”, notably, Qiu et al. (2024) and Zhong et al. (2023). Qiu et al. (2024) is motivated by testing the ability of LLMs to perform human-like induction reasoning, and Zhong et al. (2023) aims to support open-ended exploration. While similar in spirit, we examine the case of generating theories between input and labels for challenging problems where even researchers may struggle with proposing new hypotheses.
To summarize, we highlight the following contributions:
  • We propose a novel computational framework for generating and evaluating hypotheses with large language models.
  • Our generated hypotheses enable interpretable hypothesis-based classifiers that outperform in-context learning and even supervised learning for one synthetic and three real-world datasets. These hypotheses are also robust across different LLMs and out-of-distribution datasets.
  • We find our generated hypotheses to corroborate existing findings while also providing new insights for the tasks.
2 Method
We begin with a description of the problem formulation. Given a set $\mathcal{S}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ where $x_{i}$ is an example and $y_{i}$ is the corresponding label, the goal is to learn a set of hypotheses =$\mathcal{H}=\{h_{1},...,h_{m}\}$ that describe theories of relationships between $x$ and $y$. To this end, we prompt an LLM to summarize demonstration examples into high-level hypotheses (§ 2.1). Then, during inference, the LLM makes inference based on the generated hypothesis (§ 2.2). Our code is available at https://github.com/ChicagoHAI/hypothesis_generation.
2.1 Hypothesis Generation
Our hypothesis generation algorithm (Algorithm 1) is inspired by the upper confidence bound (UCB) algorithm (Auer, 2002). Given a set of initial examples $\mathcal{S}_{\mathrm{init}}\subset\mathcal{S}$, we first prompt an LLM to generate hypotheses for $\mathcal{S}_{\mathrm{init}}$, which serve as our initial hypothesis bank $\mathcal{H}$. While initialized hypotheses may explain some portions of data, they often fall short of encompassing the full scope of the examples. We thus introduce an update stage which serves a dual purpose: 1) it increases the percentage of data explainable by the hypotheses and 2) it replaces any hypotheses that are found to be inaccurate.
In the update stage, for a training example $s$, we select the top $k$ high-reward hypotheses from the hypothesis bank $\mathcal{H}$. The LLM is prompted to make a prediction with each of the top $k$ high-reward hypotheses on $s$. Then we compute the accuracy of the inference and accordingly update the reward for each of the hypotheses. If $w_{hyp}$ hypotheses predict incorrectly for the example $s$, then $s$ is added to a wrong example pool $\mathcal{W}$. Once the wrong example pool reaches a max size of $w_{max}$, the wrong examples in $\mathcal{W}$ are used to generate new hypotheses, as shown in Algorithm 1. The wrong example pool represents the gap in knowledge that the current pool of hypotheses has for the dataset. Thus by generating new hypotheses, the algorithm fills in these gaps. We update $\mathcal{H}$ with the newly generated hypotheses according to the rewards.
Reward. As mentioned above, each hypothesis has an associated reward. In our algorithm, we use the reward function in the UCB algorithm due to similarities between the multi-arm bandit problem and our problem formulation. In particular, we consider each hypothesis to be an arm and each training example to be a “pull”. We note, however, that unlike the multi-arm bandit problem, multiple hypotheses are tested for a singular train example. Moreover, there can be new arms after hypotheses are updated, altering the setting from the standard static arms scenario to a dynamic arms scenario. Formally, the reward is defined as
$r_{i}=\frac{\sum_{(x_{j},y_{j})\in\mathcal{S}_{i}}I(y_{i}=\hat{y}_{i})}{\mathcal{S}_{i}}+\alpha\sqrt{\frac{\log{t}}{|\mathcal{S}_{i}|}},$ (1)
where $\mathcal{S}_{i}$ is the set of examples that have been used to evaluate the hypothesis $h_{i}$, $t$ is train time step, and $\alpha$ is a hyperparameter that controls the exploration term. The first term in the reward function denotes the accuracy of the hypothesis for all $\mathcal{S}_{i}$. The second term is the exploration term, which is computed based on the number of times the hypothesis has been selected and the number of training examples visited so far. The accuracy term urges the algorithm to use well-performing hypotheses, whereas the exploration term encourages the algorithm to explore hypotheses that have not been selected many times. Thus, the reward function strikes a balance between exploration and exploitation.
For more details on implementation of HypoGeniC, refer to § B.1.
Algorithm 1 HypoGeniC
1:Training samples $\mathcal{S},\mathtt{num\_init},k,w_{max},H$
$\mathcal{H}$leftarrow generate_hypotheses($\{\mathcal{S}_{i}:i\leq\mathtt{num\_init}\}$) $\triangleright$ Initialize hypothesis bank
3:$\mathcal{W}$leftarrow {}
4:for $(x_{t},y_{t})\in\mathcal{S}$ :
5:   $\mathcal{H}_{\mathrm{top}}=\leftarrow\{h:\text{$h\in\mathcal{H}$ is amongst top $k$ based on reward}\}$
6:   for $h\in\mathcal{H}_{\mathrm{top}}$ :
[…]
8:      update_reward($h,y_{t},\hat{y}_{t}^{h}$)
9:   if |$|\{\mathbf{wrong}(\hat{y}_{t}^{h}):h\in\mathcal{H}\}|\geq w_{hyp}$ : $\triangleright$ $w_{hyp}$ is dynamically determined, see § B.1
10:      $\mathcal{W}$leftarrow $\mathcal{W}$$\cup\{(x_{t},y_{t})\}$
11:   if $|\mathcal{W}|=w_{max}$ :
12:      $\mathcal{N}\leftarrow$ generate_hypotheses($\mathcal{W}$)
13:      $\mathcal{W}$leftarrow {}
14:      $\mathcal{H}\leftarrow\{h:\text{$h$ is amongst top $H$ based on reward in $\mathcal{H}\cup\mathcal{N}$}\}$
15:return $\mathcal{H}$
2.2 Hypothesis-based Inference
For efficiency purposes, we use each hypothesis on its own without accounting for their combinatorial effect during training; however, we should leverage the set of hypotheses as a whole during inference for at least two reasons. Firstly, some hypotheses may only apply to a subset of examples. Second, competing theories may require head-to-head comparisons. Hence, we develop multiple inference strategies to account for these different styles of reasoning (see Appendix A for prompts and § B.2 for implementation details).
  • Best-accuracy hypothesis. The hypothesis $h$ with the highest accuracy from the hypothesis bank $\mathcal{H}$ is included in the prompt to guide the model to perform inference.
  • Filter and weighted vote. One hypothesis may not be enough to explain the data. Thus, this approach uses a combination of relevant hypotheses to make predictions for a single example. For each example, we first filter hypotheses by prompting an LLM to judge which hypotheses are relevant to the example. Next, an LLM is prompted to generate predictions for each of the relevant hypotheses, and these predictions are aggregated with weighted vote, where the weight is the training accuracy of the corresponding hypothesis.
  • Single-step adaptive inference. Similar to filter and weighted vote, this approach leverages contextual information to choose hypotheses. The difference, however, is that it selects the most applicable hypothesis for each test example. Specifically, for a given test example, the LLM is tasked with identifying the most applicable hypothesis from a set of options. For each hypothesis, we provide instances from the training set where the hypothesis was accurate. Then, the LLM selects the most relevant hypothesis by comparing the test example to these training examples and evaluating their similarity. Thereafter, we apply the hypothesis to the test example to perform inference. Please note that this is all done in one step with a long prompt.
  • Two-step adaptive inference. We divide the previous inference strategy into two steps:
    1. 1.
      The LLM determines the most relevant set of examples by comparing the test example with the corresponding examples of the hypotheses.
    2. 2.
      Then, the corresponding hypothesis is provided to the LLM, which it uses to perform inference on the test example in a second prompt.
3 Experiment Setup
Next, we introduce the experiment setup to evaluate HypoGeniC.
3.1 Tasks and Datasets
The choice of appropriate tasks is critical for evaluating the ability of LLMs to generate hypothesis. The focus of our work is on generating hypotheses based on observed data. A prerequisite is that potential hypotheses do exist. In the context of classification, it implies that the classification performance is non-trivial. In addition, we need to ensure that the hypotheses describing the data are likely not a priori known by LLMs, which rules out standard tasks such as sentiment analysis. Therefore, we use four datasets that satisfy these requirements: a synthetic task where we know the true hypothesis and three real-world datasets that exhibit complex underlying patterns and constitute widely studied social science problems.
Shoe sales is a synthetic task we created to investigate the scenario where there is only one single valid hypothesis. The task is to predict the color of the shoe that the customer will buy based on their appearance. The input provides appearance features, namely, age, height, gender, color of the hat, color of the shirt, color of the bag, and size of the bag. We construct this dataset such that the color of the shoe must match the color of the shirt. Since there are six colors in total, this becomes a 6-class classification problem.
Deceptive review detection is an instance of deception detection, a widely studied phenomenon in psychology and other social sciences (Granhag & Vrij, 2005). This particular task (Deceptive reviews) requires distinguishing genuine reviews from fictitious ones (Ott et al., 2011), where human performance is about chance (Lai & Tan, 2019). The dataset includes 800 genuine reviews and 800 fictitious reviews for 20 hotels in Chicago.
Predicting popularity is a notoriously challenging task in social sciences because it is known to be affected by seemingly random factors (Salganik et al., 2006). We use two datasets in this work: nd Tweet popularity. s derived from a dataset in the Upworthy Research Archive (Matias et al., 2021). The original dataset was collected through A/B testing, where each user was shown pairs of a headline and image for multiple packages (articles). Each user was exposed to only one of these pairs per package, and the clicks were recorded for each pair per package. [^1] The Upworthy Research Archive only provides the image IDs instead of the graphics. We thus only use the headlines for our dataset. This process resulted in a total of 150,816 headlines across 22,666 packages. We construct a binary classification dataset by choosing for each package the headline that received the most clicks and the one that received the fewest. We remove all sets of duplicate headlines, which results in our version of the ataset. The task for this dataset is to deduce which headline had more clicks in a pair. Tweet popularity uses a dataset of 13,174 tweet pairs (Tan et al., 2014), which are matched by the topic and the author. Similar to the task is to predict which one received more retweets.
3.2 Baselines, Oracles, and Evaluation Metrics
We use three different LLMs in our experiments (Mixtral (Mistral, 2023), GPT-3.5-turbo (OpenAI, 2023a), and Claude-2.1 (Anthropic, 2023)). We compare our approach with the following methods.
  1. 1.
    Zero-shot and few-shot prompting. We provide LLMs with task-specific instructions (zero-shot), optionally accompanied by three demonstration examples (few-shot).
  2. 2.
    No updates. To assess the value of the update stage in our algorithm, we evaluate the performance of the initialized hypotheses. In particular, we pick the best-performing hypothesis on the training set and use it for inference on the test set.
  3. 3.
    Supervised Learning. We fine-tune RoBERTa on each of the datasets to serve as a non-interpretable oracle. We include results for training on 200 examples and 1000 examples. Since fine-tuning update model weights, we expect RoBERTa to set the upper bound on in-distribution datasets.
We randomly sample 200 training examples and 300 test examples for each dataset. Since all our datasets are classification tasks with ground truth labels, we use accuracy as our evaluation metric. To understand the effect of the number of training examples, we evaluate the performance of all methods at 10, 25, 50, 100, and 200 training examples. We also experiment with two different hypothesis bank sizes: 3 and 20 hypotheses to evaluate the impact of utilizing a larger number of hypotheses. The detailed hyperparameters of our approach can be found in § B.3.
4 Results
To demonstrate the effectiveness of our hypothesis generation approach, we present results via three evaluation methods. First, we show that in the standard supervised learning setup, our generated hypotheses enable more accurate predictions than baselines and even oracles when using a small set of examples. Second, we evaluate the generated hypotheses by checking whether they can generalize across different inference LLMs and to out-of-distribution datasets. We find surprisingly consistent performance even when using a different LLM to make inference from the generated hypotheses. Finally, we conduct a qualitative analysis to show that the generated hypotheses not only corroborate with existing theories but also provide novel insights about the tasks at hand.
Shoe Deceptive Headline Tweet
Models Methods Sales Reviews Popularity Popularity
RoBERTa (Oracle) Train 200 100.0 84.0 49.0 50.7
Train 1000 100.0 91.0 60.0 62.0
Claude-2.1 Zero shot 36.0 31.0 59.0 50.3
Few shot 75.0 51.0 60.0 0.3*
HypoGeniC (no updates) 100.0 70.3 57.3 59.0
HypoGeniC 100.0 75.3 61.3 62.0
Mixtral Zero shot 43.0 55.0 55.0 2.7*
Few shot 79.0 56.3 55.3 48.7
HypoGeniC (no updates) 96.0 60.3 59.7 60.7
HypoGeniC 98.0 68.0 60.3 62.7
GPT-3.5-turbo Zero shot 39.0 50.0 56.0 41.0
Few shot 49.0 55.0 60.0 62.0
HypoGeniC (no updates) 100.0 56.0 44.0 45.0
HypoGeniC 100.0 60.7 63.7 61.0
Table 1: Prediction accuracies with 200 examples. We report the best numbers across all hyperparameter configurations, number of training examples, and inference strategies for HypoGeniC (we discuss their effect in details in § 4.1). The sensitive nature of the Tweet popularity dataset may cause models to have their safety mode triggered. These results are marked by * in the table.
4.1 Performance on Heldout Test Sets
As discussed in the introduction, a side product of our approach is an interpretable hypothesis-based classifier. We compare its performance with standard supervised learning with fine-tuned RoBERTa and few-shot prompt learning (Table 1).
Our generated hypotheses improve inference over standard zero-shot and few-shot inference. Across all LLMs, HypoGeniC outperforms the zero-shot learning by an average of 60% on Shoe sales, 22.7% on Deceptive reviews, 5.1% on and 30.6% on Tweet popularity. Similarly, we find that HypoGeniC shows an increase from few-shot learning by 31.7% on Shoe sales, 13.9% on Deceptive reviews, 3.3% on and 24.9% on Tweet popularity. Note that these results are inflated on Tweet popularity as safety mode is triggered for Mixtral and Claude-2.1 for zero-shot and few-shot learning respectively. The results show that hypothesis-based inference can increase the performance of LLMs significantly. One exception is that our method performs slightly worse (by 1%) than the few-shot baseline in the Tweet popularity with GPT-3.5-turbo. One possible reason is that the few-shot demonstrations are effective at eliciting the pretraining knowledge in GPT-3.5-turbo, possibly due to a large amount of tweets in pretraining data. For more detailed results, refer to Appendix C.
We also evaluate generated hypotheses with oracle inference, where the model retrospectively picks the best hypothesis for each prediction from the bank. With oracle inference, HypoGeniC achieves on average 88.6% on Deceptive reviews, 84.1% on and 88% on Tweet popularity across all LLMs, which are superior to results in Table 1. This result further suggests that hypotheses generated by HypoGeniC are of high quality and can lead to accurate predictions when the correct hypothesis is selected.
HypoGeniC matches or even exceeds RoBERTa with the same number of training examples on most datasets. Both HypoGeniC and RoBERTa yield 100% on the syntheic dataset. Moreover, HypoGeniC is 12.8% and 11.2% better than RoBERTa on nd Tweet popularity respectively with 200 training examples. Since RoBERTa learns by updating model weights to minimize the cross-entropy loss, it tends to benefit from more training examples, so we increase training examples to 1000 for RoBERTa. Despite the accuracy boost from more training examples, we find that HypoGeniC’s best result still outperforms RoBERTa by 3.7% on nd 0.7% on Tweet popularity. One exception, however, is the Deceptive reviews dataset. We suspect that as word-level features are very useful in this dataset (Ott et al., 2011), they could be tougher for LLMs to extract but easier for fine-tuned models to grasp.
Updating hypothesis bank leads to hypotheses of higher quality. Comparing HypoGeniC with the “no updates” results, we find that updating hypotheses generally leads to better hypotheses, suggesting that our algorithm is effective at improving hypothesis quality. The improvement is on average 0.7% on Shoe sales, 5.8% on Deceptive reviews, 8.1% on and 7% on Tweet popularity. Another advantage of HypoGeniC over “no updates” is that sometimes the training examples exceed the context window size of LLMs, which can lead to degraded performance (Figures 4 and 3).
Effect of inference strategy. Figure 2 shows HypoGeniC results with different inference strategies on Deceptive reviews. Single-step adaptive inference is the most effective. Generally, we find that hypotheses to be one-sided, focusing on either characteristics of truthful or deceptive reviews. We thus need to consider more than one hypothesis to make a correct prediction, so best-accuracy hypothesis or two-step adaptive inference would not be ideal for this dataset. On the other datasets, we find that the effect of inference strategy is much smaller. Best-accuracy hypothesis is sufficient for Shoe sales and and filter and weighted vote works best for Tweet popularity. Results for all datasets are in § C.1. Whichever inference strategy we use, the trend of HypoGeniC against few-shot learning and RoBERTa remains largely the same.
Generally, having more training examples and a larger hypothesis pool improves performance. We show performance for different methods as number of training examples increase in Figures 4, 3 and 5. We find HypoGeniC accuracy steadily increases as training size increases on Shoe sales, suggesting that an LLM is more likely to generate the best hypothesis given more examples. For the real-world datasets, however, the performance sometimes peaks at training size at 25 or 100 before reaching to 200. We suspect that the evaluation of the hypothesis bank would be less stable for the real-world datasets, since more than one correct hypotheses are needed for the task. We also find that using a hypothesis pool of size 20 leads to better performance than using a pool of size 3.
Figure 2: HypoGeniC results with different inference strategies on Deceptive reviews. Single-step adaptive hypothesis-based inference is generally the most effective on this dataset.
4.2 Generalization of the Generated Hypotheses
Our primary interest lies in the quality of the hypotheses. A good hypothesis should enable accurate inference by any AI model or even human and also generalize to unseen out-of-distribution dataset. In this subsection, we mix and match different LLMs for generation and inference. We also evaluate the hypotheses in deceptive review prediction on a new out-of-distribution (OOD) dataset (Li et al., 2013).
Shoe Deceptive Headline Tweet
Generation Model Inference Methods Sales Reviews Popularity Popularity
Claude-2.1 Claude-2.1 100.0 67.3 57.7 62.0
Mixtral 94.0 65.0 57.7 59.3
GPT-3.5-turbo 100.0 60.7 56.3 57.7
Mixtral Claude-2.1 99.0 69.7 59.0 58.7
Mixtral 98.0 61.3 57.7 59.3
GPT-3.5-turbo 90.0 56.7 55.3 53.0
GPT-3.5-turbo Claude-2.1 100.0 75.3 60.3 59.0
Mixtral 98.0 62.0 60.0 62.3
GPT-3.5-turbo 100.0 57.3 58.7 56.3
Table 2: Performance of cross-model generation and inference with train size = 200 using best-accuracy hypothesis inference and the best hypothesis bank size between 3 and 20.
We find that the hypotheses generated by HypoGeniC generalize across models (Table 2). Generally, we find Claude-2.1 and Mixtral to be better at inference. Thus, substituting the inference model with them lead to better performance for hypothesis generated with GPT-3.5-turbo. Subsituting Claude-2.1 and Mixtral as each other’s inference model lead to small changes in performance. On Shoe sales, the performance remains high (>90%) regardless of the inference model used.
Models OOD
RoBERTa (Oracle) 73.0 ($\downarrow$11.011.011.011.0)
Claude-2.1 Few shot 41.7 ($\downarrow$9.39.39.39.3)
Claude-2.1 HypoGeniC 74.7 ($\uparrow$4.74.74.74.7)
Mixtral Few shot 49.0 ($\downarrow$7.37.37.37.3)
Mixtral HypoGeniC 64.7 ($\uparrow$1.71.71.71.7)
GPT-3.5-turbo Few shot 52.0 ($\downarrow$3.03.03.03.0)
GPT-3.5-turbo HypoGeniC 60.7 ($\uparrow$3.43.43.43.4)
Table 3: Performance on out-of-distribution deceptive reviews.
Performance even increases for Deceptive reviews and hen using Claude-2.1 as the inference model. For the cases where performance drops from Claude-2.1 to Mixtral, the decrease is marginal: 2.3% on Deceptive reviews and 2.7% on Tweet popularity.
These results suggest that the hypotheses generated by HypoGeniC are generalizable across different LLMs, which somewhat contracts the claim in Qiu et al. (2024) that LLMs cannot reliably interpret the hypotheses. We suspect that the reason may be that our tasks only rely on natural language, while their tasks rely on notions of worlds and can be fed into symbolic interpreters.
Our generated hypotheses generalize to an out-of-distribution dataset.
Table 3 presents an overview for the OOD deceptive review dataset. This dataset differs from Deceptive reviews by including reviews from four cities sourced from different websites (Li et al., 2013). We find that HypoGeniC outperforms few-shot learning by an average of 19.1%. Despite the distribution shift, HypoGeniC surprisingly increases accuracy from Deceptive reviews by an average of 3.3%, suggesting our hypotheses generalize well to this OOD dataset. Claude-2.1 remains the best performing model. In comparison, the performance of RoBERTa drops by 11%. As a result, HypoGeniC with Claude-2.1 outperforms RoBERTa by 1.7%, demonstrating the robustness of hypothesis-based inference. Refer to § C.3 for more details.
4.3 Qualitative Analysis
For the synthetic dataset, all models are able to find the true underlying hypothesis for Shoe sales: “customers tend to buy shoes that match the color of their shirt.” For the real-world datasets, we compare our hypotheses with findings from the literature. We confirm the validity of some of our hypotheses and discover new insights about the tasks that previous studies did not touch upon. We show a few examples in Table 4, and the full list of hypotheses can be found in Appendix D.
Our hypotheses confirm useful features in existing literature. For Deceptive reviews, we find that deceptive reviews are more likely to be emotional, use superlatives, or contain information that could not have been directly experienced. Similar findings are also found by previous studies on Deceptive reviews (Lai et al., 2020; Anderson & Simester, 2014; Ott et al., 2011; Li et al., 2014). For Tweet popularity, we discover that tweets that are short and concise, with specific or relevant hashtags, or with emotional tones are more likely to be retweeted more, aligning with prior studies (Tan et al., 2014; Gligorić et al., 2019). For we find that revealing something new or using vivid language and imagery can drive engagement from readers to click on headlines. Previous studies also find these rules apply to online news headlines (Banerjee & Urminsky, 2021; Sadoski et al., 2000).
We also discover new insights with our generated hypotheses. For the Deceptive reviews dataset, truthful reviews could mention the reviewer’s purpose for staying at the hotel (e.g., business trip, vacation), but deceptive ones tend not to have this information. For we find that headlines that frame the content in a personal or relatable way are clicked more. For Tweet popularity, tweets that mention influential individuals or organizations are more likely to be retweeted.
Intriguingly, one of our hypotheses contradicts a feature engineering result. Ott et al. (2011) find that the token “future” is associated with deceptive reviews, while one of our hypotheses says that mentions of “past experiences or future travel plans” are indicative of truthfulness. This discrepancy is interesting, because the context for the token “future” is unclear. It could be mentioned in the context of future plans but could also be mentioned as a complaint about “never going to stay at the hotel in the future.” Feature engineering is limited due to the contextual ambiguity, whereas our generated hypotheses and their interpretation by LLMs overcome such limitations.
Our automatic evaluation of hypothesis quality also reflects negative findings. Given mixed evidence from previous literature on the effect of “reading ease” on headline clicks, Banerjee & Urminsky (2021) finds that reading ease negatively impacts click-through rates in hrough careful feature engineering. Consistent with this result, we found that the hypotheses that claim “straightforward” and “clear” writing to be indicative of higher click-through rates have relatively lower accuracies during training.
Dataset Finding Supported/Novel
Deceptive reviews
Deceptive reviews contain more emotional terms.
Li et al. (2014)
Truthful reviews would mention weddings or special occasions.
Using vivid language and imagery helps.
Banerjee & Urminsky (2021)
Headlines that frame the content in a personal or relatable way are clicked more.
Tweet popularity
Tweets with emotional tones are retweeted more.
Tan et al. (2014)
Mentioning influential individuals or organizations leads to more retweets.
Table 4: Selected examples of generated hypotheses (on the real-world datasets) and whether they support existing findings or are novel.
5 Additional Related Work
Concept/pattern discovery. In addition to Qiu et al. (2024) and Zhong et al. (2023) discussed in the introduction, other studies have worked along similar lines (Wang et al., 2023b; Singh et al., 2023; Piriyakulkij & Ellis, 2024). For example, similar to Qiu et al. (2024), Tenenbaum et al. (2011) is motivated by human inductive reasoning and examines concept induction in synthetic settings. Ellis et al. (2020) further learns to program concepts. Similar to Zhong et al. (2023), Pham et al. (2024) generate and refine a list of topics to achieve interpretable topic modeling for open-ended exploration. Relatedly, Honovich et al. (2022) explore the deduction of task description from examples. Our work, in contrast, focuses on hypothesis generation between the input and the label for real-world challenging tasks and uses a UCB-style reward to propose novel algorithms.
Reasoning with LLMs. Although it is not our primary goal, our results show that hypothesis-based classifiers can outperform few-shot prompting. As hypotheses may be viewed as a form of reasoning, it is related to reasoning with LLMs (Wei et al., 2022; Wang et al., 2023a, i.a.). In particular, our work differs from chain-of-thought reasoning because no predefined reasoning structure is available. Moreover, an important distinction between reasoning and hypothesis generation is that the former leverages established reasoning, while the latter requires both proposition and verification of the hypotheses, with the goal of discovering unknown knowledge.
LLMs for (social) sciences. Increasing attention has been brought to the use of LLMs in social science research (Ziems et al., 2024; Kim & Lee, 2023, i.a.). Our experiments demonstrate the potential of LLMs in generating hypotheses for social science research to discover unknown knowledge in the data. Furthermore, our approach can be extended to natural sciences for general scientific discovery.
6 Conclusion
In this work, we propose HypoGeniC, a novel method that leverages LLMs to generate hypotheses with the goal of discovering unknown knowledge. With HypoGeniC, we are not only are able to generate human-interpretable hypotheses but also achieve better predictive performance against competitive baselines and even oracles. Furthermore, our method can generalize well with different models and datasets, including open models. Notably, with our generated hypotheses, we uncover new insights in real-world tasks that are widely studied in social sciences. HypoGeniC can be directly applied to natural language related tasks in social sciences. We encourage future work to explore hypothesis generation that requires additional modalities and/or leverages existing literature.
7 Limitations
In this paper, we aim to provide a robust framework for hypothesis generation, as opposed to focusing on the optimization of results. Thus, we did not perform an extensive hyperparameter search with the generation portion of HypoGeniC. We did not adjust the value of $k$, which determines $\mathcal{H}_{\mathrm{top}}$ in Algorithm 1 to maintain efficiency. Additionatlly, we only considered the effect of using a hypothesis bank size of $3$ and $20$ to only test using an extremely small hypothesis bank size and a large one. The ideal hypothesis bank size may require further investigation. Finally, we only tested the size of our wrong example bank $w_{max}$ as $10$ to strike a balance between context window sizes and generation of good quality hypotheses. We believe that a more thorough hyperparameter search could improve the performance of our methodology.
Additionally, HypoGeniC has high latency, specifically when using inference methods that require multiple prompts. For example, the filter and weighted vote inference policy requires iterating through the top hypotheses to determine relevance and then performing inference if it is relevant. For single-step adaptive inference and best accuracy hypothesis, however, HypoGeniC is efficient. We also note that the main bottleneck with performing inference lies with performing API calls to the LLM, which is a limitation not directly related to HypoGeniC. Given that we request reasoning for all inference prompts, the procedure can be time-consuming.
Appendix A Prompts
We follow the general prompt engineering guide from Claude (Anthropic, 2023) to craft the prompts. Specifically for all the prompts we use for LLMs, we split them into instruction and user prompts. In the instruction prompt, we first set a tone and context, followed by an explicit task description, and then specify the answer format. The user prompt then includes useful information such as past examples and learned hypothesis. By the end of the user prompt, we ask the LLM to make a prediction. At generation time, we input the instruction prompt to LLMs as system prompt, wrapped by the corresponding system prompt tokens for each model. Below are some example templates for the prompts associated with each task.
A.1 Shoe Sales
Instruction Prompt
Youre a helpful assistant. Your task is given as follows:
Given a set of observations, we want to generate hypotheses that are useful for predicting the color of the shoes given the appearance of the person.
Please be concise and keep the hypotheses to be one-sentence long.
Please generate them in the format of
{1. [hypothesis].
2. [hypothesis].
<num_hypotheses>. [hypothesis].}
Only propose <num_hypotheses> possible hypotheses in total.
No need to explain the hypotheses.
User Prompt
We made some observations:
··· more examples here ···
Based on the above observations, generate <num_hypotheses> hypotheses.
Please be concise and keep the hypotheses to be one-sentence long.
Please generate them in the format of
{1. [hypothesis].
2. [hypothesis].
<num_hypotheses>. [hypothesis].}
Only propose <num_hypotheses> possible hypotheses in total.
Example 1: Hypothesis Generation.
Instruction Prompt
You are a shoe salesman and want to recommend shoes to customers. There are white, red, orange, green, blue, and black shoes.
From past experiences, you learned some patterns.
Now, at each time, you should apply the learned pattern, given below, to a new customer and recommend a shoe color.
Give an answer for the shoe color recommendation. The answer should be one color word. It has to be one of white, red, orange, green, blue, and black.
User Prompt
Our learned pattern: <hypothesis_high_reward>
New customer: <appearance> is buying a pair of shoes, the shoes should be which color?
Answer:
Example 2: Hypothesis-based Inference.
Instruction Prompt
You are a shoe salesman and want to recommend shoes to customers. There are white, red, orange, green, blue, and black shoes.
Give your answer for the shoe color recommendation. The answer should be one color word. It has to be one of white, red, orange, green, blue, and black. If you do not have enough information to make a recommendation, you should give the answer "unknown".
Give your final answer in the format of "Final answer: [answer]."
User Prompt
Here are some examples of customers with certain features buying certain products:
··· more examples here ···
New customer: <appearance> is buying a pair of shoes, the shoes should be which color?
Answer:
Example 3: Zero/Few-shot Inference.
Instruction Prompt
You are a shoe salesman and want to recommend shoes to customers. There are white, red, orange, green, blue, and black shoes.
From past experiences, you learned some patterns.
For each pattern, you will also see a couple of examples that worked for each pattern.
Choose a pattern. To do this, look at the examples of each pattern, and see which of the examples the current customer is closest to. Choose the pattern corresponding to that example.
Give an answer for the shoe color recommendation.
The answer should be one word. It has to be one of white, red, orange, green, blue, and black.
Give your final answer in the following format:
Reasoning for choosing pattern: reason,
Chosen pattern: pattern,
Reasoning for choice of prediction: reason,
Final Answer: answer
User Prompt
Here are some previously generated patterns with some example where it predicted correcly what color of shoe the customer bought.
<adaptive_info_prompt>
New customer: <appearance> is buying a pair of shoes, the shoes should be which color?
Answer:
Example 4: Example-based Hypothesis Selection and Inference. <adaptive_info_prompt> consists of several hypotheses and the corresponding examples they got correct during generation time.
A.2 Deceptive Reviews
Instruction Prompt
Youre a professional hotel review analyst.
Given a set of hotel reviews, we want to generate hypotheses that are useful for predicting whether a review is truthful or deceptive. In other words, we want to know whether the review is written by a someone who actually lived in the hotel.
Using the given examples, please propose <num_hypotheses> possible hypothesis pairs.
These hypotheses should identify specific patterns that occur across the provided reviews.
Each hypothesis should contain a pair of the following:
1. A hypothesis about what makes reviews more likely to be truthful
2. The opposite hypothesis about what makes reviews more likely to be deceptive
Generate them in the format of 1. [hypothesis], 2. [hypothesis], ... <num_hypotheses>. [hypothesis].
The hypotheses should analyze what kind of reviews are likely to be truthful or deceptive.
User Prompt
We have seen some hotel reviews:
··· more examples here ···
Please generate hypotheses that are useful for predicting whether a review is truthful or deceptive.
Propose <num_hypotheses> possible hypotheses. Generate them in the format of 1. [hypothesis], 2. [hypothesis], ... <num_hypotheses>. [hypothesis].
Proposed hypotheses:
Example 5: Hypothesis Generation.
Instruction Prompt
You are a professional deceptive detection agent and your job is to determine whether a hotel review is truthful or deceptive.
In other words, we want to know whether the review is written by someone who had real experiences with the hotel.
From past experiences, you learned a pattern.
You need to determine whether each of the patterns holds for the current hotel review, and also predict whether the current hotel review is truthful or deceptive.
Give an answer. The answer should be one word (truthful or deceptive).
Give your final answer in the format of {Final answer: answer}
User Prompt
Our learned pattern: <hypothesis_high_reward>
A hotel review is the following: <review>
Given the pattern you learned above, give an answer of whether the hotel review above is deceptive or truthful.
Think step by step.
First step: Think about which pattern can be applied to the hotel review.
Second step: Based on the pattern, is this hotel review deceptive or truthful?
Example 6: Hypothesis-based Inference.
Instruction Prompt
You are a deceptive detection agent and want to determine whether a hotel review is truthful or deceptive.
In other words, we want to know whether the review is written by a someone who actually lived in the hotel.
You need to determine whether this pattern holds for the current hotel review, and also predict whether the current hotel review is truthful or deceptive.
Give an answer. The answer should be one word (truthful or deceptive).
User Prompt
We have seen some hotel reviews:
··· more examples here ···
A hotel review is the following: <review>
Is this hotel review truthful or deceptive?
Answer:
Example 7: Zero/Few-shot Inference.
Instruction Prompt
You are a professional hotel review analyst and you are able to determine whether a hotel review is deceptive or truthful.
In other words, your job is to analyze if a hotel review review is written by someone who had genuine experiences with the hotel.
From past experiences, you learned some patterns.
For each pattern, you will also see a couple of examples that worked for each pattern.
First step: take a careful look at the examples associated with each pattern, and see which set of examples the current hotel review is most similar with. Choose and repeat the pattern corresponding to that examples set.
Next, apply the pattern on the new sample to determine whether the new hotel review is deceptive or truthful.
Finally, give an answer. The answer should be one word (deceptive or truthful).
Please give your final answer in the following format:
Reasoning for choosing pattern: reason,
Chosen pattern: pattern,
Reasoning for choice of prediction: reason,
Final Answer: answer
User Prompt
Here are some previously generated patterns with some example where it predicted correctly if a hotel review is deceptive or truthful.
<adaptive_info_prompt>
A hotel review is the following: <review>
Is this hotel review truthful or deceptive?
Think step-by-step.
Step 1: Look at the new hotel review and compare it with the set of examples associated with each provided pattern.
Step 2: Find the set of examples that is the most similar to the new hotel review, pick and repeat the pattern associated with that set of examples.
Step 3: Apply the pattern you picked to the new hotel review and predict whether the new hotel review is deceptive or truthful.
Step 4: Give your final answer.
Answer:
Example 8: Example-based Hypothesis Selection and Inference. <adaptive_info_prompt> consists of several hypotheses and the corresponding examples they got correct during generation time.
A.3 Headlines With More Clicks
Instruction Prompt
You are a professional writer for an online newspaper company.
Given a pair of headlines created for the same article, you are asked to determine which will get more clicks. It is likely that the pair of headlines shares similarities, so please focus on their differences.
What difference in two headlines leads to more clicks on one than the other?
You will be given a set of observations of the format:
Headline 1: [headline]
Headline 2: [headline]
Observation: [observation].
Based on the observations, please generate hypotheses that are useful for explaining why one headline out of the pair gets more clicked than the other.
These hypotheses should identify patterns, phrases, wordings etc. that occur across the provided examples. They should also be generalizable to new instances.
Please propose <num_hypotheses> possible hypotheses and generate them in the format of 1. [hypothesis], 2. [hypothesis], ... <num_hypotheses>. [hypothesis].
User Prompt
Here are the observations:
··· more examples here ···
Please generate hypotheses that can help determine which headlines have more clicks.
Please propose <num_hypotheses> possible hypotheses.
Generate them in the format of 1. [hypothesis], 2. [hypothesis], ... <num_hypotheses>. [hypothesis].
Proposed hypotheses:
Example 9: Hypothesis Generation.
Instruction Prompt
You are a professional writer for an online newspaper company.
Given a pair of headlines created for the same article, you are asked to determine which will get more clicks. It is likely that the pair of headlines shares similarities, so please focus on their differences.
From past experiences, you learned some patterns.
Now, at each time, you should apply the learned pattern to a new pair of headlines that are created for a new article and determine which headline gets clicked more.
The answer for the higher clicks should be in the form "Headline _" where _ is either 1 or 2.
Please give your final answer in the format of {Final Answer: Headline _.}
User Prompt
Learned pattern: <hypothesis_high_reward>
Given the pattern you learned above, predict which of the following headlines will get more clicks:
Headline 1: <headline_1>
Headline 2: <headline_2>
Think step by step.
Step 1: Think about whether the pattern can be applied to the headlines.
Step 2: Analyze the difference between "Headline 1" and "Headline 2".
Step 3: Based on the pattern, which headline is likely to get more clicks?
Example 10: Hypothesis-based Inference.
Instruction Prompt
YYou are a writer for an online newspaper company. So you are excellent at determining which headlines are more likely to cause users to click on the article.
You will be given two headlines, and determine which headline was clicked more often.
You are only to give your answer.
The answer for the higher clicks should be of the form "Headline _" where _ is either 1 or 2.
Give your final answer in the following format:
"Answer: Headline _"
User Prompt
Here are some previous examples to help you:
··· more examples here ···
Which of the following headlines has more clicks:
Headline 1: <headline_1>
Headline 2: <headline_2>
Example 11: Zero/Few-shot Inference.
Instruction Prompt
You are a professional writer for an online newspaper company.
You are excellent at determining which headlines are more likely to be clicked by users.
From past experiences, you learned some patterns.
For each pattern, you will also see a couple of examples that worked for each pattern.
Please choose a pattern. To do this, look at the examples associated with each pattern, and find which set of the examples are closest to the given pair of headlines.
Please choose the pattern corresponding to that set of examples.
The answer for the higher clicks should be of the form "Headline _" where _ is either 1 or 2.
Please give your final answer in the following format:
Reasoning for choosing pattern: reason,
Chosen pattern: pattern,
Reasoning for choice of prediction: reason,
Final Answer: answer
User Prompt
Here are some previously generated patterns with some examples where it predicted which one of the pair of headlines got more clicks.
<adaptive_info_prompt>
Which one out of the following pair of headlines will get more clicks?
Headline 1: <headline_1>
Headline 2: <headline_2>
Think step by step.
Step 1: Look at the new pair of headlines and compare them with the examples associated with each pattern.
Step 2: Find the set of examples that is closest to the given pair of headlines, and pick the pattern associated with that set of examples.
Step 3: Apply the picked pattern to the new pair of headlines. Based on that pattern, think about which one out of the pair of headlines will get more clicks.
Step 4: Give your final answer.
Example 12: Example-based Hypothesis Selection and Inference. <adaptive_info_prompt> consists of several hypotheses and the corresponding examples they got correct during generation time.
A.4 Retweeted More
Instruction Prompt
You are a social media expert. You are an expert at determining which tweet will be retweeted more.
Given a set of observations, you want to generation hypotheses that will help predict which tweet out of a pair of tweets is more likely to be retweeted.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets in each pair.
Please propose <num_hypotheses> possible hypotheses.
Please generate them in the format of 1. [hypothesis], 2. [hypothesis], ... <num_hypotheses>. [hypothesis].
Please make the hypotheses general enough to be applicable to new observations.
User Prompt
We made some observations:
··· more examples here ···
Generate hypotheses that are useful for predicting which tweet out of a pair of tweets is more likely to be retweeted.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets in each pair.
Please propose <num_hypotheses> possible hypotheses.
Please generate them in the format of 1. [hypothesis], 2. [hypothesis], ... <num_hypotheses>. [hypothesis].
Proposed hypotheses:
Example 13: Hypothesis Generation.
Instruction Prompt
You are a social media expert.
Given a pair of tweets, you are asked to predict which tweet will be retweeted more.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets.
From past experiences, you learned a pattern.
Now, at each time, you should apply a learned pattern to a pair of tweets and determine which one will get more retweets.
The answer for the higher retweets should be of the form "the _ tweet" where _ is either first or second.
Please give your final answer in the format of {Final answer: the _ tweet}
User Prompt
Our learned pattern: <hypothesis_high_reward>
The first tweet: <first_tweet>
The second tweet: <second_tweet>
Given the pattern you learned above, predict which one of the two tweets will get more retweets.
Think step by step.
First step: Think about if the pattern can be applied to the tweets.
Second step: Analyze the textual difference between the two tweets.
Third step: Based on the pattern, which tweet is more likely to get more retweets?
Final step: Give your final answer in the format of {Final answer: the _ tweet}
Final answer:
Example 14: Hypothesis-based Inference.
Instruction Prompt
You are a social media expert.
Given a pair of tweets, you are asked to predict which tweet will be retweeted more.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets.
The answer for the higher retweets should be of the form "the _ tweet" where _ is either first or second.
Please give your final answer in the format of {Final answer: the _ tweet}
User Prompt
Here are some examples:
··· more examples here ···
The first tweet: <first_tweet>
The second tweet: <second_tweet>
Which one of the two tweets will get more retweets?
Example 15: Zero/Few-shot Inference.
Instruction Prompt
You are a social media expert.
Given a pair of tweets, you are asked to predict which tweet will be retweeted more.
Please note that the paired tweets are about the same content and are posted by the same user, so you should focus on the wording difference between the two tweets.
From past experiences, you learned some patterns.
You should apply a learned pattern to a pair of tweets and determine which one will get more retweets.
For each pattern, you will also see a couple of examples that worked for each pattern.
Please choose a pattern. To do this, look at the examples associated with each pattern, and find which set of the examples are closest to the given pair of tweets.
Please choose the pattern corresponding to that set of examples.
Please give your final answer in the following format:
Reasoning for choosing pattern: reason,
Chosen pattern: pattern,
Reasoning for choice of prediction: reason,
Final Answer: answer
User Prompt
Here are some previously generated patterns with some examples where it predicted which tweet will will be retweeted more.
<adaptive_info_prompt>
The first tweet: <first_tweet>
The second tweet: <second_tweet>
Which one of the two tweets will get more retweets?
Think step by step.
Step 1: Look at the new pair of tweets and compare them with the examples associated with each pattern.
Step 2: Find the set of examples that is closest to the given pair of tweets, and pick the pattern associated with that set of examples.
Step 3: Analyze the textual difference between the two tweets.
Step 4: Apply the picked pattern to the new pair of tweets. Based on that pattern, think about which one out of the pair of headlines will get more clicks.
Step 5: Give your final answer.
Example 16: Example-based Hypothesis Selection and Inference. <adaptive_info_prompt> consists of several hypotheses and the corresponding examples they got correct during generation time.
Appendix B Implementation Details
B.1 HypoGeniC implementation
Sampling
When initializing the rewards of newly generated hypotheses, we use the examples in the wrong example bank to do so. Given that we work in a low data regime, for hypotheses generated near the end of the training loop, the accuracies of hypotheses are likely to be biased. To counter this phenomenon, we also allow for the hypotheses to use the initial examples $\mathcal{S}_{\mathrm{init}}$ for initializing rewards. By allowing the hypotheses to initialize reward with more examples, the accuracy lies closer to its true value, allowing for fair comparison between earlier generated hypotheses and newer ones.
Dynamic hypotheses update
In Algorithm 1, we display how we generate and update the hypotheses pool $\mathcal{H}$. In particular, we add an example $s$ to the wrong example bank $\mathcal{W}$ if the number of hypotheses that incorrectly predict $s$ is greater than $w_{hyp}$. In our implementation, we use a linearly increasing $w_{hyp}$ as training time $t$ increases. This allows our algorithm to update the hypotheses more frequently at early stage of training, and less frequently at the end.
B.2 Inference method implementations
Filter and weighted vote
In order to filter the hypotheses, we iterate through the top $k$ hypotheses ranked by reward. For each hypothesis, we ask the Large Language Model (LLM) if it is relevant. Thereafter, for each of the relevant hypotheses, the LLM is prompted to use the hypothesis to make predictions. Then, for each predicted label, we add up the accuracy scores from the hypotheses that outputted that particular label. The final label is the one that has highest total accuracy score.
One-step adaptive and two-step adaptive inference
The detailed framework of our adaptive inference methods is split into two parts - hypotheses pruning and hypotheses selection. In the case where we have a large number of hypotheses, it is likely that some hypotheses in $\mathcal{H}$ have overlaps or are paraphrases of each other.
We address this issue with the following procedure:
  1. 1.
    During training, we record the examples that each hypothesis correctly predicts.
  2. 2.
    Then we create one-hot encodings for each hypothesis, where the $i$-th element of the one-hot encoding is 1 if the hypothesis correctly predicts the $i$-th example, and 0 otherwise. We subsequently compute a similarity matrix between each pair of hypotheses by taking the pairwise cosine similarities.
  3. 3.
    Lastly, we create a linear program with the objective of maximizing the sum of accuracies of the selected hypotheses, subject to the constraint that every pair of the selected hypotheses has a similarity score below a predefined threshold $\gamma$.
After pruning the set of hypotheses, we prompt the LLM to pick one hypothesis for its final prediction, as described in § 2.2. For the single-step adaptive inference, we ask the LLM to select a hypothesis and make a prediction in one prompt. On the other hand, with the two-step adaptive inference, we first prompt the LLM to select a hypothesis and then prompt the LLM again to make a prediction based on the selected hypothesis.
B.3 Hyperparameters
For the training stage, we set a limit on the hypothesis bank size, experimenting with sizes $H=3$ and $H=20$ to determine the impact of utilizing a larger number of hypotheses. Throughout all the experiments, we use the reward coefficient $\alpha=0.5$, $w_{max}=10$, $\mathtt{num\_init}=10$, and we have two different sets of the rest of hyperparameters for hypothesis bank sizes of $3$ and $20$.
  • With $H=3$, we use $k=2$ and generate $1$ hypothesis per update. For inference, we employ all $3$ hypotheses for filter and weighted vote. For single-step and two-step adaptive inference, we use all $3$ hypotheses with $\gamma=0.3$ and provide $5$ examples to each hypothesis.
  • In the case of $H=20$, we use $k=10$ and generate $5$ hypotheses per update. Then we take the top $5$ hypotheses, ranked by their training accuracies, for filter and weighted vote. For single-step and two-step adaptive inference, we use the top $5$ hypotheses with $\gamma=0.7$ and provide $5$ examples each.
Appendix C Detailed Results
C.1 HypoGeniC Performance across inference strategies
(a) Shoe sales
(b) Deceptive reviews
(e) HypoGeniC results with different inference strategies. Best-accuracy hypothesis is sufficient for getting good performance on Shoe sales and Single-step adaptive hypothesis-based inference is the most effective on Deceptive reviews. Filter and weighted vote is best on Tweet popularity.
(d) Tweet popularity
§ C.1 presents the best results for all of our inference strategies, considering every dataset and all hyperparameter configurations.
For Shoe sales, we observe that all the models perform effectively by using the best hypothesis inference strategy. Surprisingly, Mixtral is unable to perform perfectly. This is because despite generating the hypothesis that fully describes the data, Mixtral opts not to apply the hypotheses, favoring to choose a random label for the sake of “variety”. Both GPT-3.5-turbo and Mixtral display similar patterns across the inference strategies, with best-accuracy hypothesis, filter and weighted vote, and two-step adaptive inference all having comparable performance. However, for all models we find single-step adaptive inference drops in accuracy. Given that two-step adaptive inference performs well, it is likely that the long prompt causes the model difficulty in choosing the correct hypotheses. For Claude-2.1, we see that filter and weighted vote drops in performance. As this method searches for relevant hypotheses, the model is likely finding that inaccurate patterns relevant, which end up outweighing the inference of the best hypothesis.
For Deceptive reviews, Claude-2.1 is the best performing model across all inference policies. Across the models, we highlight that single-step adaptive inference method works best for this dataset. In this inference method, the prompt specifically includes the aims of determining if a review is deceptive. This likely helps the model use the context provided to better decide which set of example resembles the test example most. Hence, splitting up the prompt may have caused performance to suffer.
We find that s the most challenging dataset. As mentioned in § 3.1, the original dataset was created with both images and headlines paired together. In our version of the dataset, we only use the headlines, so we are missing a crucial variable that contributes to understanding click behavior. Therefore, based off only headlines, it is difficult to generate hypotheses that truly capture the data. Despite this challenge, we note that our hypotheses can still adeptly capture a large portion of data with 63.7% being our highest accuracy. Specifically, we find that the best-accuracy hypothesis strategy performs best. We also note that filter and weighted vote can provide strong performance as in the case of Claude-2.1 and GPT-3.5-turbo, suggesting that hypotheses corroborating with each other can lead to better performance. We observe that GPT-3.5-turbo is the best performing model here, with all inference policies (aside from single-step adaptive) having high accuracy.
Finally, over the Tweet popularity dataset, we find that the filter and weighted vote is the best choice for inference policy, with it being the best inference method for GPT-3.5-turbo and Mixtral. This indicates that using hypotheses in conjunction is useful as multiple variables together adeptly characterize the dataset. The performance of the rest of the inference policies has no clear pattern over this dataset.
(e) HypoGeniC results with different inference strategies. Best-accuracy hypothesis is sufficient for getting good performance on Shoe sales and Single-step adaptive hypothesis-based inference is the most effective on Deceptive reviews. Filter and weighted vote is best on Tweet popularity.